The table below gives the list price and the number of bids received for five randomly selected items sold through online auctions. Using this data, consider the equation of the regression line, yˆ=b0+b1x, for predicting the number of bids an item will receive based on the list price. Keep in mind, the correlation coefficient may or may not be statistically significant for the data given. Remember, in practice, it would not be appropriate to use the regression line to make a prediction if the correlation coefficient is not statistically significant. Price in Dollars 21 37 40 41 44 Number of Bids 2 5 6 7 10 Table Step 1 of 7 : Find the estimated slope. Round your answer to three decimal places. 1. Find the estimated slope. Round your answer to three decimal places. 2.Find the value of the coefficient of determination. Round your answer to three decimal places. 3.Find the estimated y-intercept. Round your answer to three decimal places 4.Find the estimated value of y when x=44. Round your answer to three decimal places.Step 4 of 6: Determine the value of the dependent variable y^ at x=0. Step 5 of 6: Substitute the values you found in steps 1 and 2 into the equation for the regression line to find the estimated linear model. According to this model, if the value of the independent variable is increased by one unit, then find the change in the dependent variable y^.
Answer #1

Data:

Price in $ No. of bids
21 2
37 5
40 6
41 7
44 10

No. of bids vs Price($) 12 10 y = 0.2916X - 4.6731 R? = 0.8234 8 6 4 N 0 0 10 20 30 40 50

Equation: No. of bids = 0.2916 Price($) - 4.673

1. Find the estimated slope. Round your answer to three decimal places.

Slope = 0.292

2.Find the value of the coefficient of determination. Round your answer to three decimal places.

R-sq = 0.823

3.Find the estimated y-intercept. Round your answer to three decimal places

Intercept = -4.673

4.Find the estimated value of y when x=44. Round your answer to three decimal places.

No. of bids = -4.673 + 0.292*44 = 8.175

Step 4 of 6: Determine the value of the dependent variable y^ at x=0.

No. of bids = -4.673 + 0 = -4.673 (negative)

Step 5 of 6: Substitute the values you found in steps 1 and 2 into the equation for the regression line to find the estimated linear model.

substititing new values:

Data:

Price in $ No. of bids
21 2
37 5
40 6
41 7
44 10
44 8.175
0 -4.673

No. of bids vs Price($) 12 10 y = 0.2917x -4.6746 R2 = 0.9578 00 6 4 N 0 0 10 20 30 40 50 -2 -4 -6

According to this model, if the value of the independent variable is increased by one unit, then find the change in the dependent variable y^.

When value of x increases by one unit, value of y increases by 0.292 units.

Please rate my answer and comment for doubt